OpenStudy (anonymous):

HELP ME PLEASE !!!! What is the rule for the function shown in the table? X|-1|0|1|2 Y|-2|1|4|7

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

A) y= 1/3 x + 1 B) y= 3x + 1 C) y= 1/3x D) y= 1/2 x +1